OpenStudy (anonymous):

solve: (x^(4/3)*y^(2/3)/(xy)^(1/2)

OpenStudy (anonymous):

answer: y^1/6 * x^5/6

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

I don't know how to solve it though

OpenStudy (anonymous):

oooh i thought you posted the solution someone else did

OpenStudy (anonymous):

subtract the exponents \[\large x^{\frac{4}{3}-\frac{1}{2}}\] and \[\large y^{\frac{2}{3}-\frac{1}{2}}\]

OpenStudy (anonymous):

when you are dividing with the same base, subtract the exponents so for x the exponent will be (4/3 - 1/2), that's how you get the xraised to 5/6 same with your y

OpenStudy (anonymous):

wait a second the last part should be 1/3 not 1/2

OpenStudy (anonymous):

my answer is xy^(1/3) is that right?